The length of a rectangle is 10 feet longer than it is wide. If each side is increased 10 feet, then the area is multiplied by 2. What was the size of the original rectangle?

The width or short side is ___ feet

The length or long side is ___ feet
Answer #1

Let width, w=a. Then length, L = 10+x. Area of erect angle, A= WXL= xx (104x) it each side is increased, w=10+2 4-(1072)+10 L